OpenStudy (anonymous):

Solve the inequality: 2< or equal to 1-1/4x

OpenStudy (mertsj):

\[2\le1-\frac{1}{4}x\]

OpenStudy (mertsj):

Subtract 1 from both sides.

OpenStudy (mertsj):

\[1\le-\frac{1}{4}x\]

OpenStudy (mertsj):

Multiply both sided by -4 and don't forget to reverse the inequality sign.

OpenStudy (anonymous):

so x>=.25?

OpenStudy (anonymous):

or negative .25??

OpenStudy (mertsj):

\[-4\ge x\]

OpenStudy (anonymous):

Ohh see my calculator said 1/4 times 1 is .25 I get it

OpenStudy (mertsj):